The length of a rectangle is 16 centimeters less than five times its width. Its area is 45 square centimeters. Find the dimensions of the rectangle.
Answer by addingup(3677)

L x W = 45
L = 5W-16 substitute for L in the equation above:
5W-16 x W = 45 distribute the W (Distributive Properties of Multiplication) and subtract 45 on both sides:
5W^2-16W-45 = 0
Quadratic formula:
W = [-(-16)+-sqrt((-16)^2-4(5)(-45))]/2(5)
W = (16+-sqrt1156)/10
W = (16+-34)/10
W = 50/10 or W = 18/10
W = 5 or W = -1.8
Since we are not looking for a negative number, let's try the 5:
L = 5W-16 = 5(5)-16 = 25-16 = 9
L x W = 45
9 x 5 = Correct, you now have the correct answer.